# Graph Attention-Driven Bayesian Deep Unrolling for Dual-Peak Single-Photon Lidar Imaging

### v1

@@ I Introduction
In this paper, we propose an algorithm unrolling method for dual-peak single-photon Lidar imaging. We extend an existing Bayesian model [32] and its iterative algorithm for dual-peak imaging. Each step of the iterative method is unrolled into neural network layers. Unlike existing deep learning approaches that adopt image-based representations, our method employs a point-cloud representation for supporting dual peaks and utilizes graph neural networks (GNNs) for feature extraction. By combining the strengths of statistical modeling with the adaptability of deep learning, our architecture is both efficient and interpretable, providing uncertainty information on the inferred point clouds. An early version of the proposed work was presented in [37]. We extend this conference paper to provide a new Bayesian model considering multiple peaks, improve network architecture, and add uncertainty estimation.
Processing large histogram data requires a large amount of memory and high computational cost. To address this issue, we adopt a multiscale approach [32], where we downsample the histogram data and estimate the dual peaks from each downsampled histogram. The estimated dual peaks are used as initial multiscale point clouds, which is the input to the proposed network. In this regard, we compressed large histogram data [38, 39] into a multiscale point cloud representation, which is more efficient to process and requires less memory. The proposed network layers sequentially estimate the true depth from the multiscale point cloud and refine the multiscale point cloud. The components of layers rely on soft attention and hard attention inspired by median filtering in the Bayesian algorithm [32]. The proposed method is evaluated on synthetic and real data and shows superior performance compared to existing methods with uncertainty information, a key advantage of the proposed method.
@@ II Multiscale multi-peak observation model
where $r_{n,k}$ and $d_{n,k}$ are the reflectivity and the depth value of the $k$-th target, respectively, $b_{n}$ is the background photons from ambient light or detector noises and $g$ is the system impulse response function (IRF). This IRF is commonly approximated by the Gaussian function $\mathcal{N}(t\,;\mu,\sigma^{2})$ with the mean $\mu$ and the variance $\sigma^{2}$, satisfying $\sum_{t=1}^{T}g\left(t-d_{n,k}\right)=1$ [6, 32]. With the assumption of independent observations with respect to $n$ and $t$, the likelihood function of the observation model is given by
$p\,(\boldsymbol{Y}\mid\boldsymbol{d},\boldsymbol{r},\boldsymbol{b})=\prod_{n=1}^{N}\prod_{t=1}^{T}\frac{s_{n,t}^{y_{n,t}}}{y_{n,t}!}\exp^{-s_{n,t}},$ (2)
where $\boldsymbol{Y}=\{y_{n,t}\}$, $\boldsymbol{d}=\{d_{n,k}\}$, $\boldsymbol{r}=\{r_{n,k}\}$, and $\boldsymbol{b}=\{b_{n}\}$. This likelihood function (2) for the $n$-th pixel can be rewritten as
@@
×$\times\prod_{k}\exp\left(-r_{n,k}\sum_{t}g(t-d_{n,k})\right).$
@@
×$\times\prod_{t,k}g(t-d_{n,k})^{y_{n,t}},$
where $Q$ is a function of $\boldsymbol{y}_{n}$ and $\bar{r}_{n}:=\sum_{k=1}^{K}r_{n,k}$ approximating the total photon counts in the $n$-th pixel $\sum_{t=1}^{T}y_{n,t}$. The approximated lower bound of the likelihood function can then be rewritten as
@@
×$\times\overline{Q}\left(\boldsymbol{y}_{n}\right),$
@@
×$\times\overline{Q}\left(\boldsymbol{y}_{n}^{(\ell)}\right).$
@@ III-A Prior and posterior distributions
We have derived the likelihood function $p\,(\boldsymbol{Y}\mid\boldsymbol{D})$ in (7) and the prior distribution $p(\boldsymbol{x},\boldsymbol{D}\mid\boldsymbol{\epsilon},\boldsymbol{W})$ in (8) and $p(\boldsymbol{\epsilon})$ in (9), which leads to the posterior distribution proportional to
@@ III-B Iterative algorithm
[…]
This operation will be replaced by attention mechanisms in the proposed deep learning model in Section IV.
Expansion step. The multiscale depths $\boldsymbol{d}^{(1,\cdots,L)}$ appear in the likelihood (5) and the prior (8). Considering minimizing the negative logarithm of the posterior distribution, the update of the multiscale depths is given by
[…]
@@
7:  ab Squeeze step: Update $\boldsymbol{x}$
@@
Figure 1: Overview of the proposed network architecture for one stage. GAT stands for Graph Attention Networks and kNN stands for k-Nearest Neighbor method.
@@ IV-A Network
To learn meaningful features from the multiscale point cloud, we construct a graph by applying the k-nearest neighbors (kNN) algorithm with $k=6$. On the constructed graph, we use Graph Attention Networks (GATs) [41] to extract features from the point cloud and compute attention weights. The GAT generalizes the convolutional operation on the graph, allowing the network to learn the importance of each neighboring point.
@@ IV-B Estimation of Initial Multiscale Dual Peaks
[…]
@@ V-A Results on synthetic data
Comparison methods. We compare the proposed method with other single-photon Lidar imaging methods by Lindell et al. [25], Peng et al. [27], and BU3D [31]. As these state-of-the-art methods consider one surface per pixel, they are not directly applicable to the dual-peak reconstruction. We run these methods twice to estimate the depth of the first and the second surfaces separately, assuming the interval of each surface. In the first run, the histogram data is modified to eliminate the second surface, copying uniform background photons onto the time bins between 400 and 700. In the second run, the histogram is modified to eliminate the first surface, copying uniform background photons onto the time bins between 0 and 300. Note that, when running the proposed method, we do not modify the histogram data and directly estimate the dual peaks.
@@
PPP $=16.0$, SBR$=4.0$
Refer to caption Refer to caption Refer to caption Refer to caption
PPP $=4.0$, SBR$=4.0$
Refer to caption Refer to caption Refer to caption Refer to caption
PPP $=4.0$, SBR$=1.0$
Refer to caption Refer to caption Refer to caption Refer to caption
Lindell [25] (2 runs) Peng [27] (2 runs) BU3D [31] (2 runs) Proposed
Figure 3: Reconstructed point clouds on the Art scene with dual peaks. The first three columns show the results by Lindell, Peng and BU3D, respectively, with two runs, assuming that the approximate positions of surfaces are known. The last column shows the results by the proposed method.
@@
PPP $=16.0$, SBR$=4.0$
Refer to caption Refer to caption Refer to caption Refer to caption
PPP $=4.0$, SBR$=4.0$
Refer to caption Refer to caption Refer to caption Refer to caption
PPP $=4.0$, SBR$=1.0$
Refer to caption Refer to caption Refer to caption Refer to caption
Lindell [25] (2 runs) Peng [27] (2 runs) BU3D [31] (2 runs) Proposed
Figure 5: Reconstructed point clouds on the Bowling scene with dual peaks. The first three columns show the results by Lindell, Peng and BU3D, respectively, with two runs, assuming that the approximate positions of surfaces are known. The last column shows the results by the proposed method.
@@
TABLE I: Quantitative comparison on the Art scene. Chamfer distance consists of two terms: O-To-G (distance from the estimated point cloud to the ground truth) and G-To-O (distance from the ground truth to the estimated point cloud).
@@
DAE Chamfer dist. O-To-G G-To-O DAE Chamfer dist. O-To-G G-To-O DAE Chamfer dist. O-To-G G-To-O
@@
Lindell (2 runs) 0.0261 0.0130 0.0099 0.0031 0.0131 0.0087 0.0060 0.0027 0.0122 0.0082 0.0055 0.0026
Peng (2 runs) 0.0099 0.0079 0.0054 0.0024 0.0082 0.0070 0.0049 0.0021 0.0078 0.0068 0.0047 0.0021
BU3D (2 runs) 0.0275 0.0112 0.0078 0.0035 0.0116 0.0077 0.0044 0.0033 0.0098 0.0073 0.0040 0.0033
Proposed 0.0274 0.0180 0.0157 0.0023 0.0093 0.0059 0.0041 0.0018 0.0074 0.0045 0.0028 0.0017
@@
Lindell (2 runs) 0.0120 0.0079 0.0053 0.0026 0.0140 0.0089 0.0061 0.0028 0.0153 0.0097 0.0067 0.0030
Peng (2 runs) 0.0063 0.0060 0.0041 0.0018 0.0053 0.0055 0.0037 0.0017 0.0049 0.0052 0.0035 0.0017
BU3D (2 runs) 0.0086 0.0072 0.0041 0.0032 0.0064 0.0064 0.0034 0.0030 0.0061 0.0062 0.0034 0.0029
Proposed 0.0058 0.0033 0.0021 0.0012 0.0047 0.0023 0.0013 0.0010 0.0045 0.0021 0.0012 0.0009
@@
Lindell (2 runs) 0.0268 0.0171 0.0121 0.0050 0.0299 0.0203 0.0140 0.0063 0.0305 0.0211 0.0144 0.0067
Peng (2 runs) 0.0039 0.0044 0.0029 0.0015 0.0033 0.0040 0.0026 0.0015 0.0031 0.0039 0.0025 0.0015
BU3D (2 runs) 0.0054 0.0057 0.0032 0.0025 0.0044 0.0050 0.0028 0.0022 0.0042 0.0047 0.0026 0.0021
Proposed 0.0039 0.0014 0.0007 0.0006 0.0037 0.0011 0.0006 0.0005 0.0036 0.0010 0.0005 0.0005
TABLE II: Quantitative comparison on the Bowling scene. Chamfer distance consists of two terms: O-To-G (distance from the estimated point cloud to the ground truth) and G-To-O (distance from the ground truth to the estimated point cloud).
@@
DAE Chamfer dist. O-To-G G-To-O DAE Chamfer dist. O-To-G G-To-O DAE Chamfer dist. O-To-G G-To-O
@@
Lindell (2 runs) 0.0183 0.0108 0.0078 0.0030 0.0095 0.0075 0.0048 0.0026 0.0086 0.0070 0.0045 0.0025
Peng (2 runs) 0.0075 0.0064 0.0043 0.0021 0.0066 0.0059 0.0041 0.0019 0.0065 0.0058 0.0040 0.0019
BU3D (2 runs) 0.0162 0.0085 0.0054 0.0031 0.0067 0.0065 0.0035 0.0030 0.0061 0.0064 0.0033 0.0030
Proposed 0.0188 0.0115 0.0095 0.0020 0.0062 0.0045 0.0003 0.0016 0.0051 0.0038 0.0023 0.0015
@@
Lindell (2 runs) 0.0081 0.0066 0.0042 0.0024 0.0097 0.0075 0.0050 0.0025 0.0109 0.0082 0.0055 0.0026
Peng (2 runs) 0.0053 0.0051 0.0034 0.0017 0.0046 0.0047 0.0031 0.0016 0.0043 0.0045 0.0029 0.0016
BU3D (2 runs) 0.0056 0.0063 0.0033 0.0030 0.0049 0.0060 0.0031 0.0029 0.0049 0.0059 0.0031 0.0028
Proposed 0.0035 0.0026 0.0015 0.0011 0.0028 0.0019 0.0011 0.0008 0.0027 0.0017 0.0010 0.0008
@@
Lindell (2 runs) 0.0238 0.0155 0.0111 0.0044 0.0273 0.0185 0.0126 0.0058 0.0283 0.0194 0.0130 0.0063
Peng (2 runs) 0.0034 0.0039 0.0025 0.0014 0.0030 0.0036 0.0022 0.0014 0.0029 0.0036 0.0022 0.0014
BU3D (2 runs) 0.0043 0.0053 0.0029 0.0024 0.0037 0.0046 0.0026 0.0021 0.0035 0.0044 0.0024 0.0020
Proposed 0.0021 0.0012 0.0006 0.0005 0.0019 0.0009 0.0005 0.0004 0.0018 0.0008 0.0004 0.0004
@@
Qualitative comparison. Fig. 3 reports the reconstructed point clouds for the Art scene. Note that existing deep learning methods of Lindell, Peng and BU3D are run twice to estimate the depth of the first and the second surfaces separately. The results by Peng observe bleeding artifacts, while the result by BU3D show less such artifacts, but observe some outlier points in the third row. Compared to other methods, the proposed method reconstructs the point clouds with less bleeding artifacts with less outliers, but often has outliers in the far region in the third row. Fig. 4 shows the estimated depth maps for the same scene. Overall, the results are consistent with the point cloud results. The last column shows the uncertainty information by the proposed method. The uncertainty map shows higher values in the regions with outliers, indicating the uncertainty of the depth estimation.

### v2

@@ I Introduction
In this paper, we propose an algorithm unrolling method for dual-peak single-photon Lidar imaging. We extend an existing Bayesian model [32] and its iterative algorithm for dual-peak imaging. Each step of the iterative method is unrolled into neural network layers. Unlike existing deep learning approaches that adopt image-based representations, our method employs a point-cloud representation for supporting dual peaks and utilizes graph neural networks (GNNs) for feature extraction. By combining the strengths of statistical modeling with the adaptability of deep learning, our architecture is both efficient and interpretable, providing uncertainty information on the inferred point clouds. An early version of the proposed work was presented in [37]. We extend this conference paper to provide a new Bayesian model considering multiple peaks, add uncertainty estimation, and improve the network architecture by incorporating an explicit hard attention mechanism and modifying the graph attention layer design.
Processing large histogram data requires a large amount of memory and high computational cost. To address this issue, we adopt a multiscale approach [32], where we downsample the histogram data and estimate the dual peaks from each downsampled histogram. The estimated dual peaks are used as initial multiscale point clouds, which is the input to the proposed network. In this regard, we compressed large histogram data [38, 39] into a multiscale point cloud representation, which is more efficient to process and requires less memory. The proposed network layers sequentially estimate the true depth from the multiscale point cloud and refine the multiscale point cloud. The components of layers rely on soft attention and hard attention inspired by median filtering in the Bayesian algorithm [32]. The proposed method is evaluated on synthetic and real data and shows superior performance compared to existing methods with uncertainty information, a key advantage of the proposed method. Our code is available at https://github.com/daedalus-KM/PointcloudUnrolling.
@@ II Multiscale multi-peak observation model
where $r_{n,k}$ and $d_{n,k}$ are the reflectivity and the depth value of the $k$-th target, respectively, $b_{n}$ is the background photons from ambient light or detector noises and $g$ is the system impulse response function (IRF). We assume that the background photons are absent for simplicity. This IRF is commonly approximated by the Gaussian function $\mathcal{N}(t\,;\mu,\sigma^{2})$ with the mean $\mu$ and the variance $\sigma^{2}$, satisfying $\sum_{t=1}^{T}g\left(t-d_{n,k}\right)=1$ [6, 32]. With the assumption of independent observations with respect to $n$ and $t$, the likelihood function of the observation model is given by
$p\,(\boldsymbol{Y}\mid\boldsymbol{d},\boldsymbol{r})=\prod_{n=1}^{N}\prod_{t=1}^{T}\frac{s_{n,t}^{y_{n,t}}}{y_{n,t}!}\exp^{-s_{n,t}},$ (2)
where $\boldsymbol{Y}=\{y_{n,t}\}$, $\boldsymbol{d}=\{d_{n,k}\}$ and $\boldsymbol{r}=\{r_{n,k}\}$. This likelihood function (2) for the $n$-th pixel can be rewritten as
@@
$\times\prod_{k}\exp\left(-r_{n,k}\right).$
@@
$\times\prod_{t,k}g(t-d_{n,k})^{y_{n,t}},$
where $Q$ is a function of $\boldsymbol{y}_{n}$ and $\bar{r}_{n}:=\frac{1}{K}\sum_{k=1}^{K}r_{n,k}$ approximating the scaled total photon counts in the $n$-th pixel $\frac{1}{K}\sum_{t=1}^{T}y_{n,t}$. The approximated lower bound of the likelihood function can then be rewritten as
@@
$\times\overline{Q}\left(\boldsymbol{y}_{n}\right),$
@@
$\times\overline{Q}\left(\boldsymbol{y}_{n}^{(\ell)}\right).$
@@ III-A Prior and posterior distributions
Let $\mathbf{D}=[\boldsymbol{d}^{(1)},\boldsymbol{d}^{(2)},\ldots,\boldsymbol{d}^{(L)}]$. We have derived the likelihood function $p\,(\boldsymbol{Y}\mid\boldsymbol{D})$ in (7) and the prior distribution $p(\boldsymbol{x},\boldsymbol{D}\mid\boldsymbol{\epsilon},\boldsymbol{W})$ in (8) and $p(\boldsymbol{\epsilon})$ in (9), which leads to the posterior distribution proportional to
@@ III-B Iterative algorithm
$x_{n,k}\leftarrow\underset{x}{\operatorname{argmin}}\,\mathcal{C}(x),\,\text{where }\mathcal{C}(x)=\sum_{\ell,n^{\prime}\in\nu_{n}}w_{n^{\prime},n,k}^{(\ell)}\left|x-d_{n^{\prime},k}^{(\ell)}\right|.$ (11)
This operation will be replaced by attention mechanisms in the proposed deep learning model in Section IV. Expansion step. The multiscale depths $\boldsymbol{d}^{(1,\cdots,L)}$ appear in the likelihood (5) and the prior (8). Considering minimizing the negative logarithm of the posterior distribution, the update of the multiscale depths is given by
$d_{n,k}^{(\ell)}\leftarrow\underset{d}{\operatorname{argmin}}\frac{\left[d-d_{n,k}^{\mathrm{ML}(\ell)}\right]^{2}}{2\bar{\sigma}^{2(\ell)}}+\sum_{n^{\prime}\in\nu_{n}}\frac{w_{n,n^{\prime},k}^{(\ell)}\left|d-x_{n^{\prime}}\right|}{\epsilon_{n^{\prime},k}}.$ (12)
@@
7: Squeeze step: Update $\boldsymbol{x}$
@@
Figure 1: Overview of the proposed network architecture for one stage. GAT stands for Graph Attention Networks and kNN stands for k-Nearest Neighbor.
@@ IV-A Network
To learn meaningful features from the multiscale point cloud, we construct a graph by applying the k-nearest neighbors (kNN) algorithm with $k=6$. On the constructed graph, we use Graph Attention Networks (GATs) [41] with a single head to extract features from the point cloud and compute attention weights. The GAT generalizes the convolutional operation on the graph, allowing the network to learn the importance of each neighboring point.
@@ IV-B Estimation of Initial Multiscale Dual Peaks
To estimate the initial multiscale point cloud from the original histogram data, we apply the cross-correlation to the histogram by the IRF. To the resulting histogram, we apply the convolution spatially with 4 uniform filters whose sizes are {1×1, 3×3, 7×7, $\{1\times 1,\,3\times 3$$\times 7$$\times 13\}$. This spatial convolution yields low-pass filtered histograms $\boldsymbol{Y}^{(1,2,3,4)}$. To support reconstructing two surfaces per pixel, we estimate dual peaks from each histogram $\boldsymbol{Y}^{(1,2,3,4)}$, as the input to our network is an initial multiscale point cloud rather than the large volume histogram.
@@ V-A Results on synthetic data
Comparison methods. We compare the proposed method with other state-of-the-art single-photon Lidar imaging methods by RT3D [18], Lindell et al.[25], Peng et al.[27], and BU3D [31]. While RT3D supports multi-surface reconstruction, other three methods consider one surface per pixel and are thus not directly applicable to the dual-peak reconstruction. We run these methods [25, 27, 31] twice to estimate the depth of the first and the second surfaces separately, assuming the interval of each surface. In the first run, the histogram data is modified to eliminate the second surface, copying uniform background photons onto the time bins between 400 and 700. In the second run, the histogram is modified to eliminate the first surface, copying uniform background photons onto the time bins between 0 and 300. Note that, when running the proposed method, we do not modify the histogram data and directly estimate the dual peaks.
@@
PPP =16.0=16.0, SBR=4.0=4.0
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PPP =4.0=4.0, SBR=4.0=4.0
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PPP =4.0=4.0, SBR=1.0=1.0
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
RT3D [18] Lindell [25] (2 runs) Peng [27] (2 runs) BU3D [31] (2 runs) Proposed
Figure 3: Reconstructed point clouds on the Art scene with dual peaks. The first column shows the results by RT3D. The next three columns show the results by Lindell, Peng and BU3D, respectively, with two runs, assuming that the approximate positions of surfaces are known. The last column shows the results by the proposed method.
@@
PPP =16.0=16.0, SBR=4.0=4.0
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PPP =4.0=4.0, SBR=4.0=4.0
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PPP =4.0=4.0, SBR=1.0=1.0
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
RT3D [18] Lindell [25] (2 runs) Peng [27] (2 runs) BU3D [31] (2 runs) Proposed
Figure 5: Reconstructed point clouds on the Bowling scene with dual peaks. The first column shows the results by RT3D. The next three columns show the results by Lindell, Peng and BU3D, respectively, with two runs, assuming that the approximate positions of surfaces are known. The last column shows the results by the proposed method.
@@
Figure 6: Reconstructed depth maps with different PPP and SBR levels on the Bowling scene. The first column shows the ground truth depth map with first peak (top), and second peak (bottom)
TABLE I: Quantitative comparison on the Art scene. Chamfer distance consists of two terms: O-To-G (distance from the estimated point cloud to the ground truth) and G-To-O (distance from the ground truth to the estimated point cloud). DAE, O-To-G, and G-To-O values are reported as mean $\pm$ standard deviation of pixelwise measures. All values are scaled by $10^{-2}$ for readibility.
@@
DAE Chamfer O-To-G G-To-O DAE Chamfer O-To-G G-To-O DAE Chamfer O-To-G G-To-O
@@
RT3D $1.92$ $0.88\pm 0.91$ $04\pm 0.90 1.1$71.17 $0.51\pm 0.52$ $0.66\pm 0.61$ $1.08$ $\pm 0.51 0.54±0.45$$\pm 0.45$
Lindell (2 runs) $2.61\pm 12.42$ $1.55$ $1.28\pm 3.49$ $0.27\pm 0.21$ $1.31\pm 4.21$ $0.98$ $0.75\pm 1.30$ $0.23\pm 0.19$ $1.22\pm 3.75$ $0.91$ $0.69\pm 1.13$ $0.22\pm 0.19$
Peng (2 runs) $\textbf{0.99}\pm 2.32$ 0.85 $\textbf{0.66}\pm 0.91$ $0.20\pm 0.17$ $\textbf{0.82}\pm 1.93$ $0.74$ $0.57\pm 0.82$ $0.17\pm 0.14$ $0.78\pm 1.85$ $0.71$ $0.55\pm 0.80$ $0.16\pm 0.13$
BU3D (2 runs) $2.74\pm 14.83$ $1.22$ $0.88\pm 3.72$ $0.35\pm 0.25$ $1.16\pm 6.91$ $0.80$ $0.46\pm 1.59$ $0.33\pm 0.23$ $0.98\pm 5.49$ $0.75$ $0.41\pm 1.15$ $0.33\pm 0.23$
Proposed $2.73\pm 12.52$ $1.92$ $1.73\pm 9.71$ $\textbf{0.18}\pm 0.18$ $0.93\pm 4.76$ 0.58 $\textbf{0.44}\pm 3.04$ $\textbf{0.15}\pm 0.16$ $\textbf{0.74}\pm 3.07$ 0.43 $\textbf{0.29}\pm 1.12$ $\textbf{0.14}\pm 0.15$
@@
RT3D $1.01$ $0.40\pm 0.53$ $0.61\pm 0.67$ $0.60$ $0.25\pm 0.24$ $0.35\pm 0.36$ $0.67$ $0.40\pm 0.40$ $0.26\pm 0.20$
Lindell (2 runs) $1.20\pm 3.07$ $0.92$ $0.71\pm 1.20$ $0.21\pm 0.18$ $40\pm 3.29 1.1$31.13 $0.89\pm 1.46$ $0.24\pm 0.20$ $1.53\pm 3.41$ $1.27$ $1.01\pm 1.57$ $0.26\pm 0.23$
Peng (2 runs) $0.63\pm 1.50$ $0.61$ $0.47\pm 0.72$ $0.14\pm 0.11$ $0.53\pm 1.23$ $0.54$ $0.41\pm 0.62$ $0.13\pm 0.11$ $0.49\pm 1.16$ $0.52$ $0.39\pm 0.56$ $0.13\pm 0.10$
BU3D (2 runs) $0.86\pm 4.08$ $0.74$ $0.42\pm 1.44$ $0.31\pm 0.18$ $\pm 2.07 0.640.64$ $0.35\pm 0.44$ $0.29\pm 0.17$ $0.61\pm 1.93$ $0.62$ $0.34\pm 0.37$ $0.27\pm 0.17$
Proposed $\textbf{0.58}\pm 3.11$ 0.30 $\textbf{0.20}\pm 1.71$ $\textbf{0.10}\pm 0.13$ $\textbf{0.47}\pm 2.44$ 0.20 $\textbf{0.12}\pm 0.51$ $\textbf{0.08}\pm 0.13$ $\textbf{0.45}\pm 2.28$ 0.18 $\textbf{0.11}\pm 0.15$ $\textbf{0.07}\pm 0.12$
@@
RT3D $0.71$ $0.35\pm 1.02$ $0.36\pm 0.34$ $0.53$ $0.29\pm 0.32$ $0.24\pm 0.19$ $0.50$ $9\pm 0.25 0.22±0$$2\pm 0.15$
Lindell (2 runs) $2.68\pm 3.94$ $2.51$ $2.00\pm 2.09$ $0.51\pm 0.53$ $2.99\pm 3.86$ $2.97$ $2.28\pm 1.98$ $0.69\pm 0.67$ $3.05\pm 3.79$ $3.10$ $2.33\pm 1.90$ $0.76\pm 0.72$
Peng (2 runs) $\textbf{0.39}\pm 0.97$ $0.42$ $0.30\pm 0.43$ $0.12\pm 0.10$ $\textbf{0.33}\pm 0.78$ $0.38$ $0.26\pm 0.33$ $0.11\pm 0.10$ $\textbf{0.31}\pm 0.72$ $0.37$ $0.25\pm 0.30$ $0.11\pm 0.10$
BU3D (2 runs) $0.54\pm 2.34$ $0.56$ $0.33\pm 0.80$ $0.23\pm 0.16$ $0.44\pm 1.44$ $0.47$ $0.28\pm 0.29$ $0.20\pm 0.15$ $0.42\pm 1.35$ $0.45$ $0.26\pm 0.23$ $0.18\pm 0.14$
Proposed $\textbf{0.39}\pm 2.28$ 0.12 $\textbf{0.07}\pm 0.24$ $\textbf{0.05}\pm 0.11$ $37\pm 2.24 0.0$9 $\textbf{0.05}\pm 0.10$ $\textbf{0.04}\pm 0.11$ $36\pm 2.24 0.0$8 $\textbf{0.04}\pm 0.09$ $\textbf{0.04}\pm 0.10$
TABLE II: Quantitative comparison on the Bowling scene. Chamfer distance consists of two terms: O-To-G (distance from the estimated point cloud to the ground truth) and G-To-O (distance from the ground truth to the estimated point cloud). DAE, O-To-G, and G-To-O values are reported as mean $\pm$ standard deviation of pixelwise measures. All values are scaled by $10^{-2}$ for readibility.
@@
DAE Chamfer O-To-G G-To-O DAE Chamfer O-To-G G-To-O DAE Chamfer O-To-G G-To-O
@@
RT3D $1.71$ $0.70\pm 0.68$ $1.01\pm 0.89$ $1.02$ $0.42\pm 0.44$ $0.60\pm 0.52$ $1.07$ $0.57\pm 0.53$ $0.51\pm 0.39$
Lindell (2 runs) $1.83\pm 9.44$ $1.44$ $1.17\pm 3.01$ $0.27\pm 0.20$ $\pm 2.57 0.950.95$ $0.72\pm 1.19$ $0.23\pm 0.18$ $0.86\pm 2.17$ $0.87$ $0.65\pm 1.04$ $0.22\pm 0.17$
Peng (2 runs) $\textbf{0.75}\pm 1.70$ 0.77 $\textbf{0.59}\pm 0.82$ $0.18\pm 0.15$ $0.66\pm 1.50$ $0.69$ $0.54\pm 0.78$ $0.16\pm 0.12$ $0.65\pm 1.47$ $0.68$ $0.53\pm 0.78$ $0.15\pm 0.12$
BU3D (2 runs) $1.61\pm 11.03$ $1.00$ $0.67\pm 2.86$ $0.33\pm 0.22$ $0.67\pm 3.88$ $0.72$ $9\pm 0.93 0.33±0$$3\pm 0.20$ $0.61\pm 2.92$ $0.70$ $7\pm 0.63 0.33±0$$3\pm 0.19$
Proposed $1.88\pm 9.93$ $1.39$ $1.22\pm 7.12$ $\textbf{0.16}\pm 0.16$ $\textbf{0.62}\pm 3.77$ 0.48 $\textbf{0.35}\pm 2.08$ $\textbf{0.13}\pm 0.13$ $\textbf{0.51}\pm 2.82$ 0.39 $\textbf{0.26}\pm 0.77$ $\textbf{0.12}\pm 0.13$
@@
RT3D $0.98$ $0.41\pm 0.69$ $0.56\pm 0.60$ $0.54$ $0.26\pm 0.26$ $0.28\pm 0.26$ $0.63$ $0.38\pm 0.36$ $0.25\pm 0.19$
Lindell (2 runs) $\pm 2.05 0.810.81$ $0.61\pm 1.01$ $0.21\pm 0.16$ $0.97\pm 2.20$ $0.98$ $0.76\pm 1.26$ $0.22\pm 0.16$ $09\pm 2.31 1.1$01.10 $0.86\pm 1.38$ $0.24\pm 0.17$
Peng (2 runs) $0.53\pm 1.26$ $0.57$ $0.43\pm 0.66$ $0.13\pm 0.11$ $0.46\pm 1.08$ $0.50$ $0.37\pm 0.55$ $0.13\pm 0.10$ $0.43\pm 0.99$ $0.48$ $0.35\pm 0.51$ $0.12\pm 0.10$
BU3D (2 runs) $0.55\pm 2.13$ $0.69$ $6\pm 0.77 0.33±0$$3\pm 0.16$ $0.49\pm 1.51$ $0.64$ $0.34\pm 0.30$ $0.30\pm 0.15$ $0.49\pm 1.43$ $0.63$ $0.34\pm 0.29$ $0.29\pm 0.15$
Proposed $\textbf{0.35}\pm 2.51$ 0.24 $\textbf{0.16}\pm 0.86$ $\textbf{0.09}\pm 0.11$ $\textbf{0.28}\pm 1.99$ 0.17 $\textbf{0.10}\pm 0.19$ $\textbf{0.07}\pm 0.10$ $\textbf{0.27}\pm 2.07$ 0.16 $\textbf{0.09}\pm 0.15$ $\textbf{0.06}\pm 0.10$
@@
RT3D $0.73$ $0.35\pm 1.40$ $0.38\pm 0.39$ $0.51$ $0.28\pm 0.45$ $0.23\pm 0.17$ $0.46$ $0.26\pm 0.23$ $0.20\pm 0.12$
Lindell (2 runs) $2.38\pm 2.82$ $2.50$ $2.05\pm 2.05$ $0.45\pm 0.31$ $2.73\pm 2.60$ $3.03$ $2.40\pm 1.86$ $0.63\pm 0.39$ $2.83\pm 2.50$ $3.22$ $2.50\pm 1.75$ $0.72\pm 0.43$
Peng (2 runs) $0.34\pm 0.83$ $0.39$ $0.28\pm 0.36$ $0.12\pm 0.10$ $0.30\pm 0.69$ $0.36$ $0.24\pm 0.28$ $0.11\pm 0.10$ $0.29\pm 0.65$ $0.35$ $0.24\pm 0.26$ $0.12\pm 0.10$
BU3D (2 runs) $0.43\pm 1.34$ $0.56$ $0.31\pm 0.46$ $0.24\pm 0.14$ $0.37\pm 1.07$ $0.47$ $0.27\pm 0.24$ $0.20\pm 0.13$ $0.35\pm 0.97$ $0.44$ $0.26\pm 0.22$ $0.18\pm 0.13$
Proposed $\textbf{0.21}\pm 1.90$ 0.10 $\textbf{0.06}\pm 0.14$ $\textbf{0.04}\pm 0.08$ $\textbf{0.19}\pm 1.81$ 0.07 $\textbf{0.04}\pm 0.12$ $\textbf{0.03}\pm 0.08$ $\textbf{0.18}\pm 1.78$ 0.06 $\textbf{0.03}\pm 0.09$ $\textbf{0.03}\pm 0.07$
@@
Qualitative comparison. Fig. 3 reports the reconstructed point clouds for the Art scene. Note that existing deep learning methods of Lindell, Peng and BU3D are run twice to estimate the depth of the first and the second surfaces separately. The results by Peng observe bleeding artifacts, while the results by BU3D show fewer such artifacts but observe some outlier points in the third row. Meanwhile, RT3D shows fewer bleeding artifacts as well, but yields a sparser reconstruction in the third row with fewer points. Compared to other methods, the proposed method reconstructs the point clouds with less bleeding artifacts with less outliers, but often has outliers in the far region in the third row. Fig. 4 shows the estimated depth maps for the same scene. Overall, the results are consistent with the point cloud results. The last column shows the uncertainty information by the proposed method. The uncertainty map shows higher values in the regions with outliers, indicating the uncertainty of the depth estimation.